# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

### v1

@@ 5.1.2. HOW
CataractBot provides both an immediate LLM-generated response and an asynchronous expert-verified response. For time-sensitive questions, participants reported using Google to “double-check” the initial LLM-generated, unverified response due to the inherent delay associated with obtaining a verified response. Four participants, accustomed to relying on Google for their informational needs, reserved CataractBot exclusively for questions requiring expert opinions. At times, they first used Google and subsequently posed the same question to the bot, hoping for a quick, verified response from the expert. In only two instances involving critical questions, participants relied on the bot’s response and took interim actions before getting verification from the expert due to the time-sensitive nature of the questions. According to D4, without CataractBot, patients might have resorted to Google, risking following inaccurate advice, given the unavailability of doctors for immediate consultations. However, CataractBot improves the likelihood of receiving accurate responses, as it is utilizing a custom knowledge base. Irrespective of time sensitivity, certain questions such as “What time i should come?” (P19), or “I get operated on 5th jan, can i do exercises as alom vilom etc?” (P27), could only be adequately answered by hospital experts. Participants relied on CataractBot for such questions.
We found that the experts were “fine” with the additional task of verifying the bot’s responses, as articulated by D4:
“It takes only about 15-20 seconds to answer each question… they are easy. It was not really eating into my time where I had like hundred messages [to respond] every day… It’s much easier than replying to my mailbox.” – D4.
None of the doctors adhered to a specific time for responding; they verified queries while awaiting the next patient, during lunch or coffee breaks, post-OPD towards the end of the day after 4 pm, and/or after reaching home. The only exception was during days reserved for surgeries. All doctors emphasized a no-phone-use policy during surgical procedures, emphasizing the need for minimal messages on those days. Even the coordinators suggested adding a working-time parameter for CataractBot, restricting message delivery to the specified period.
As the experts (except one coordinator) used CataractBot on their personal WhatsApp accounts, it blurred the boundary between their home and work (similar to (Mols and Pridmore, 2021; Wang et al., 2020)). On a positive note, WhatsApp minimized the learning curve for experts. However, our experts pointed out that WhatsApp had a low signal-to-noise ratio, resulting in instances where they missed CataractBot messages: “I am not much into WhatsApp. I just see messages in the important groups, that’s it. Otherwise my WhatsApp has 5000+ [unread] messages… college groups, family groups.” (D2). Addressing this issue, the escalation doctor D4 suggested that experts should be recommended to “pin” CataractBot on WhatsApp, ensuring it remains consistently at the top of their chat list. Additionally, experts managed multiple patients simultaneously through CataractBot. It could potentially compromise the quality of patient care, as observed by Johansson et al. (2020a). For instance, an expert mistakenly tagged the wrong query while adding a correction, leading to confusion for the patient.
Learning Curve: A majority of participants found the bot easy to use, with 86.2% rating ‘Agree’ or ‘Strongly Agree’ in response to the usability question. Several factors contributed to the ease of learning, with participants’ familiarity with WhatsApp being the most significant. Additionally, features such as support for multiple languages, audio support, and related questions (Figure 2), facilitated a quick start. Encouraging patients to use the bot as part of onboarding and providing detailed instructions proved essential in minimizing the learning curve. Contrary to initial expectations that understanding the three icons might be complex, participants found it self-explanatory:
“That is very simple… Even my dad understood. I didn’t tell him the tick mark, question mark… He told me, “okay, doctor verified… Now the tick mark came”… Even for the tick mark also you are sending notification. If you would have done it in the background without a message, he would have not noticed… If MY dad was able to use, I think anyone can use, because even attending a phone call is a bit difficult for him.” – A13.
Explicit notification messages related to the experts’ verification step were appreciated by multiple participants.
5.1.3. WHAT
To conduct a more detailed analysis of the medical and logistical questions posed by participants, one of the researchers manually classified them into 12 broad categories, identified in a bottom-up manner (similar to (Li et al., 2023; Clarke et al., 2016b)). Note: Several questions fell into multiple categories. Figure 4 illustrates the relative significance of these categories before, during, and after the day of surgery. Prior to the surgery, patients and attendants were concerned about the surgical procedure (asking questions about “anaesthesia” and “duration of surgery”) and sought guidance on how to prepare for it (e.g., “eat or not before the surgery”). One day before the surgery, with the coordinator finalizing the schedule, questions related to the “schedule” peaked, as patients/attendants awaited calls from patient coordinators. On the day of surgery, with patients starting a course of eye drops to prevent infection and reduce inflammation, questions about medication and dosage rose in importance, as did those about the expected discharge procedures. After the surgery, participants were particularly concerned about pain management and sought guidance on dos-and-don’ts during recovery, including information on exercise and travel.
Regarding question structure, our participants minimally used double-barreled questions—single messages containing multiple questions or topics—unlike prior work (Li et al., 2023). This difference may be attributed to demographic variations, as Li et al. (2023) focused on doctors who aimed to maximize productivity by asking multiple questions together. Only four double-barreled questions were sent by three of our participants, and in each instance, the two questions were related. E.g., “What is the Valsalva maneuver and why is it harmful after cataract surgery?” (A21). The bot answered both questions in each case.
Anthropomorphization: A total of 39 messages from 5 patients and 13 attendants showed signs of anthropomorphizing the bot. Participants often addressed the bot as ‘Doctor’, ‘Sir’, or ‘Ma’am’. Many queries were prefaced with greetings (like ‘hi’) and included acknowledgements (like ‘okay’), contributing to the humanization of the bot. In Indic languages, patients employed honorifics when addressing the bot in their messages and during interviews: “How should it be during the day? Please let me know.” – P14, in Kannada. Notably, two patients with low literacy levels treated the chat as if they were conversing back-and-forth with a human, providing non-actionable status updates, such as “The date of the surgery is tomorrow. We will leave home at 8 am from Ayyappa Nagar.” (P29). Such messages not only caused confusion for the bot but also for the experts. The classifier erred in not classifying these messages as small talk; instead, it specifically labeled them as medical to ensure that user queries were not overlooked. The observed behaviors can, in part, be attributed to participants’ awareness that doctors and patient coordinators are on the other side reviewing these messages. Participants attempted to direct their queries or provide status updates to them.
5.1.4. WHO
Patients sent a total of 89 messages (4.75.4 message/patient), while attendants sent 254 messages (8.56.8 message/attendant), significantly more than patients, with t(47)=-2.04, p¡0.05. Among the 309 verified medical and logistical answers, 62.5% were verified by the expert doctor/coordinator and the rest 37.5% by the escalation expert. This highlights the significance of the bot supporting multiple information-seeker and information-verifier roles. Further, in two cases, patients came accompanied with multiple attendants, suggesting the need for the bot to accommodate not just a patient-attendant pair, but rather, a flexible number of information-seekers per patient.
During our interviews, we identified several reasons for the patients to use the bot minimally, particularly related to demographics. Patients were typically older with lower levels of literacy and tech-literacy compared to their attendants – for example, 2 attendants, who were software testers by profession, mentioned that they “kept pushing some questions” (A21) to identify bugs and help improve the system. Moreover, a subset of patients enrolled on the bot (7) placed trust in the hospital to provide them with all the required information and did not actively seek information themselves. For instance:
“Dad didn’t try… I was the one thinking about it [the surgery]. He was like, “It’s OK! Not everyone needs to bother in the family… (or know) about what will be this, what will be that”.” – A3.
This aligns with prior findings indicating less-educated older populations may have reduced information needs (Ma et al., 2023). Additionally, patients were advised to minimize screen use post-surgery, contributing to their reduced interaction with the bot.
Patients found value in local language interaction and speech as input. A13 and A14 noted that the bot’s Indic language support helped their parents (patients) to use the bot independently.
“Without even asking me for help, he (patient) was able to use that bot as it supports Tamil… I felt it’s very useful. Because now, for elders… it is very hard to use apps and things.” – A13.
Quantitatively within patients, P13 sent the maximum number of messages to CataractBot at 16, comprising 15 text messages and 1 tap on a related question. The second-highest number of messages (15 messages) was sent by P19, the youngest patient at 41 years and a well-educated background holding a Master’s degree. Following that, P26 and P31 sent 13 and 9 messages respectively, with the majority being audio messages (8 and 6 audio messages, respectively).
Interestingly, the involvement of additional stakeholders had a negative impact, as patients tended to offload the knowledge-seeking ‘work’ to their active attendants. For instance, “I have been using it. I told him [patient] what (answers) I got. So maybe he did not feel the need (to use it).” (A20).
Moving on to experts, both the doctors and patient coordinators highlighted the significant role of escalation experts. Among the 223 answered medical questions, 76.2% were verified by operating doctors, and 23.7% by the escalation doctor. In contrast, only 26.7% of the 86 answered logistical questions were verified by the coordinator, and 73.2% by the escalation coordinator. This discrepancy occurred because the patient coordinator participated in the study using their work phone, which was only active from 9 am to 6 pm, as they left their work phone at the hospital. Moreover, they were on emergency leave for 11 days during the study period. Doctors mentioned having 3 or more days a week reserved for surgeries, and on these “operating days”, they rarely checked their phones during work hours. Despite the crucial role of escalation experts, there was a notable difference in their interactions compared to operating doctors and coordinators. While the latter had previous face-to-face meetings with patients and attendants, there was no such guarantee for escalation experts. As a result, escalation experts lacked knowledge of patients’ context, and hence provided generic responses.
5.1.5. WHEN
Out of the total 343 messages sent by patients and attendants, all, except four, were sent between 7 am and 11 pm, with an average of 14.313.7 messages/hour. Notably, the maximum number of messages (56) were sent between 10-11 am, despite participants receiving reminder messages from CataractBot at 7:30 am and 4 pm daily. The messages were distributed throughout the day, as A18 mentioned: “Whenever I am idle at work or at home… whatever (question) comes to my mind, I just ask.”.
The study was designed for patients and attendants to use the bot seven days pre-surgery and seven days post-surgery, for a maximum total of 15 days. However, due to a bug in the code to disable the bot seven days post-surgery, the bot remained active. Moreover, 3 participants had their surgeries rescheduled due to increased blood sugar levels and travel, leading them to use the bot for more than seven days before their surgery. Overall, participants sent messages to the bot from 19 days before their surgeries to 24 days post their surgeries (Figure 5). Four participants requested the bot to be available beyond the seven days post-surgery period. They cited that post-operative medications and precautions would continue for a month, and having the bot available during that time would serve as reassurance and align with the expected level of service from the hospital.
A majority of messages (71.4%) were sent in the 8-day period from 5 days prior to surgery to 2 days post-surgery. For the remaining days, across 49 users, the total number of messages was less than 10 messages/day. The highest number of messages in a single day were sent on the day of the surgery (76), and 156 messages sent before the surgery day and 111 post the surgery day. We conducted a one-way ANOVA on the number of messages asked by each patient-attendant group pre-surgery, on the day of surgery, and post-surgery, and found statistical significance with F(2, 60)=3.69, p¡.05. Post-hoc pairwise comparisons with Bonferroni correction indicated a statistically significant difference between ‘pre’ and ‘on’ surgery (t(30)=2.7, p¡.05) and between ‘pre’ and ‘post’ surgery (t(30)=3.2, p¡.05). This suggests that patients and attendants are more inquisitive before surgery, potentially due to heightened anxiety. This finding aligns with prior research indicating that patients and attendants tend to seek information closer to their doctor’s visit (Beisecker and Beisecker, 1990). The higher pre-surgery usage may also be due to the novelty effect. With respect to the types of queries, logistical questions were mainly asked before the surgery, reaching their peak one day before the scheduled surgery day (Figure 4), as anticipated. Despite the high pre-surgery usage, four participants mentioned that since they did not have access to CataractBot in the early stages of diagnosis, before scheduling the surgery, information-seeking patients and attendants had already clarified most of their questions from other sources, including Google, friends and family, and hospital staff. Particularly, three patients were scheduled for surgery in the next 2 days, so their pre-surgery usage was limited to less than 2 days.
In contrast, experts expressed the view that questions on the day of surgery might be unnecessary and suggested encouraging patients/attendants to talk directly with doctors and coordinators face-to-face:
“There are multiple questions asked on the day of surgery, “when he is going to be discharged?” Asking a question when you are in the hospital itself, looks a little too much… There is no need for you to use the bot… absolutely avoidable. You can walk 10 steps and go to the staff and ask.” – D1.
Regarding the usage patterns of experts, responses were verified throughout the day and even late at night (1 am to 2 am). The majority of responses (31) were provided by experts at 4-5 pm, post their busy workday, as most OPD and surgeries finish by 4 pm. Also, the reminder message was sent to the experts at 4 pm. Note: Every four hours during working hours (at 8 am, 12 pm, and 4 pm), a list of all questions pending verification for more than six hours was sent to both the operating and escalation experts. The second-highest responses (28) occurred at 1-2 pm, during their lunch break. This time-of-day usage pattern was described as:
@@ 5.1.5. WHEN
“I use my mobile data only when I go for breaks… lunch break, coffee break, or in the morning before I start OPD. The rest of the time, my (cellular) data is off.” – D1.
Experts refrained from using their phones in front of patients/attendants due to the negative perception associated with using phones for personal reasons. As C1 described: “If I see my phone (in front of a patient)… they will not understand what I am doing, they will think okay, simply they are watching videos or texting.”. In a busy hospital context, physical patients always took priority. Hospitals in India often face high patient loads. Experts are typically overwhelmed with more patients than they can handle. During working hours, there is usually a long queue of patients waiting to consult with experts, making it challenging for experts to use their phones comfortably to verify bot responses. This also contributed to many medical and logistical queries being escalated. We observed varying opinions on using the chatbot after work hours; two experts were reluctant to use the bot outside of working hours, while four were in favor of using it “anytime”.
@@ 5.2. Features
5.2.1. Expert-in-the-loop
Out of the 343 messages, 225 were medical queries, 87 logistical questions, 27 small talk, and 4 others. Note: 13 questions were marked as ‘Send to Coordinator’ by one of the operating doctors. On manual analysis, the escalation doctor confirmed that all of them were indeed medical questions, such as “after how many days can I shave”, but were simple enough to be answered by the patient coordinator. For our analysis, we consider them as logistical questions, as they were finally answered by the patient coordinator. Doctors approved 69.8% (157) of the bot’s generated answers directly and provided corrections for 27.1% (61) of responses, while patient coordinators approved 58.6% (51) of responses to logistical questions directly and provided corrections for 36.8% (32) of responses. For 5 medical queries and 3 logistical queries, the expert answered the poll response with a ”No” but did not provide the correction. The high number of logistical corrections was mainly because a majority of logistical questions (62.1%) were about scheduling, such as “time of surgery” and “discharge time”, for which the LLM cannot provide an appropriate response as it was not connected with the hospital’s patient management system. The length of these corrections in characters varied with 154.4105.0 for medical questions and 51.042.6 for logistical questions.
Despite having a curated knowledge base and an expanding knowledge base with the addition of expert corrections after KB Update Expert approval, the bot provided IDK responses to 16.6% of questions. Even after excluding scheduling-related questions, 23.9% of responses were corrected by the expert. Further analysis revealed that their suggested edits across all questions mainly fell into five categories: adding missing information (76.8%), asking clarification question (14.6%), factual corrections (11.0%), redirecting the patient to check with other sources (8.8%), and removing extra unnecessary information (3.7%).
@@ 5.2.1. Expert-in-the-loop
The novel feature of having human experts verify and update the auto-generated responses was highly appreciated by our participants, serving as the key reason for their engagement with the bot.
“That’s why I gave my WhatsApp (number to the coordinators while onboarding). Because I can talk directly to the doctor.” – A2.
@@
However, since participants were aware that these messages were being reviewed by doctors and coordinators, four participants used the bot to send direct messages to them instead of asking questions. For instance, P19 messaged “Give me call?”
@@
Experts did not have access to the final updated answer sent to patients/attendants after their suggested edits. We made this design decision for CataractBot to limit verification to a single round, minimizing the workload on experts. However, without this transparency, experts were sometimes unsure whether or how CataractBot had understood their corrections, especially when the edits aimed at improving readability and structure rather than solely focusing on accuracy and completeness: “I’d asked to flip [the answer] around… to give information for local anesthesia first then for general anesthesia… Did the bot actually understand and give that as an answer? I did not know.” (D3). In the CataractBot system, only the KB Update Expert had access to the updated answer sent to the information seeker.
Finally, asynchronous expert correction may not be effective in all instances, especially when patients ask critical time-sensitive questions, such as allergic reactions after surgery. In such cases, two patients suggested having a fallback option of a human agent available over a call to assist: “similar to Amazon… if the query is not resolved after two-three attempts, it should just ask you to contact [call] the coordinator.” (P21). Even the escalation doctor D4 suggested a similar solution. However, prior work (Weisz et al., 2019) suggests that in such solutions, users tend to bypass the automated agent to directly connect with the human agent over a voice call.

### v2

@@ 6.1.2. Workflow
If participants received an IDK response, they sought information from other sources, including Google and asking friends and family. For time-sensitive questions, participants reported using Google to “double-check” the initial LLM-generated, unverified response due to the delay in obtaining an expert verified response. Four participants, accustomed to relying on Google for their informational needs, reserved CataractBot exclusively for questions requiring expert opinions, such as “When can i do exercises as anulom vilom?” (P27). According to D4, without CataractBot, patients might have resorted to “Google” and risked following inaccurate advice, given the unavailability of doctors for immediate consultations. CataractBot improves the likelihood of receiving accurate responses, as it utilizes a custom KB.
We found that the experts were “fine” with the additional task of verifying the bot’s responses, as articulated by D4: “It takes only about 15-20 seconds to answer a question… It was not really eating into my time… It’s much easier than replying to my mailbox.” The majority of responses (31, 10.8%) were provided by experts at 4-5 pm, post their busy workday, as most OPD and surgeries finish by 4 pm. The second-highest responses (28, 9.8%) occurred at 1-2 pm, during their lunch break. This usage pattern was described as:
@@
“I use my mobile phone only when I go for breaks… lunch break, coffee break, or in the morning before I start OPD. The rest of the time, my (cellular) data is off.” – D1.
As five experts used CataractBot on their personal WhatsApp accounts, it minimized the learning curve but also blurred the boundary between their home and work, similar to previous findings (Mols and Pridmore, 2021; Wang et al., 2020). Our experts also noted that WhatsApp had a low signal-to-noise ratio, resulting in instances where they missed CataractBot messages: “I am not much into WhatsApp. I just see messages in the important groups. Otherwise my WhatsApp has 5000+ [unread] messages… college groups, family groups.” (D2). Addressing this issue, the escalation doctor D4 suggested that experts should be recommended to “pin” CataractBot on WhatsApp, ensuring it remains consistently at the top of their chat list.
6.1.3. Question Type and Style
To analyze the medical and logistical questions posed by participants, one of the researchers manually classified them into 12 broad categories, identified in a bottom-up manner (similar to (Li et al., 2023; Clarke et al., 2016b)). Note: Several questions fell into multiple categories. Figure 4A illustrates the relative significance of these categories before, on, and after the day of surgery. Prior to the surgery, patients and attendants were concerned about the surgical procedure (asking questions about “anaesthesia”, “lens”, and “duration of surgery”) and sought guidance on how to prepare for it (e.g., “eat or not before the surgery”, “documents to bring for surgery”, and “take blood thinners on the day of surgery”). One day before the surgery, questions related to the “schedule” peaked, as information seekers awaited calls from patient coordinators about their surgery time. On the day of surgery, questions focused on discharge procedures and post-surgery medication and dosage. After the surgery, participants were concerned about pain management and sought guidance on dos-and-don’ts during recovery (e.g., “washing hair”, “doing yoga”, and “watching TV”).
39 messages from 5 patients and 13 attendants showed signs of anthropomorphizing the bot with addressing the bot as ‘Doctor’, ‘Sir’, or ‘Ma’am’. Many queries were prefaced with greetings (like ‘hi’) and included acknowledgements (like ‘okay’), contributing to the humanization of the bot. In Indic languages, patients employed honorifics when addressing the bot in their messages and during interviews. Notably, two patients with low literacy levels treated the chat as if they were conversing back-and-forth with a human, providing non-actionable status updates, such as P29 telling CataractBot that “The date of the surgery is tomorrow. We will leave home at 8 am from Ayyappa Nagar.”
6.1.4. Multiple Stakeholders
Patients sent a total of 89 messages (4.75.4 message/patient), while attendants sent 254 messages (8.56.8 message/attendant), significantly more than patients, with t(47)=2.04, p¡0.05 (Figure 4C). Among the 309 verified medical and logistical answers, 62.5% were verified by the expert doctor/coordinator and the rest 37.5% by the escalation expert. This highlights the significance of the bot supporting multiple information seeker and information verifier roles. During our interviews, we identified a few reasons for the patients to minimally use the bot, particularly related to demographics. Patients were typically older with lower levels of literacy and tech-literacy compared to their attendants These patients placed trust in the hospital to provide them with all the required information and did not actively seek information themselves. For instance:
“Dad didn’t try… I was the one thinking about it [the surgery]. He was like, “It’s OK! Not everyone needs to bother in the family… (to know) about what will be this, what will be that”. – A3.
Additionally, patients were advised to minimize screen use post-surgery, contributing to their reduced bot interaction. Interestingly, patients sometimes offloaded the knowledge-seeking ‘work’ to their active attendants—for instance, “I have been using it. I told him [the patient] what (answers) I got. So maybe he did not feel the need (to use it).” (A20)—so the involvement of additional stakeholders had a partially negative impact..
Moving on to expert stakeholders, both the doctors and patient coordinators highlighted the significant role of escalation experts. Among the medical questions, 76.2% were verified by operating doctors and 23.7% by the escalation doctor. Doctors mentioned having 2-3 days a week reserved for surgeries, and on these “operating days”, they rarely checked their phones during work hours, leading to more escalations. Despite the crucial role of escalation experts, there was a notable difference in their interactions compared to operating doctors and coordinators. While the latter had previous face-to-face meetings with patients and attendants, there was no such guarantee for escalation experts. As a result, escalation experts lacked knowledge of patients’ context, and hence provided generic responses.
Among the different stakeholders, bot interactions varied. For patients and attendants, the bot operates as a user-driven platform where they initiate conversations by asking questions. In contrast, the bot functions as a bot-driven system for experts, initiating and driving verification-related conversations. Patients and attendants appreciated that CataractBot “does not disturb” them and provides information solely based on the questions asked. However, a few participants suggested the bot should proactively offer relevant information. For example, four attendants wished for regular updates from the chatbot during the surgery while waiting anxiously outside. A9 wanted reminders from the chatbot to ensure patients “take eye drops on time” post-surgery.
6.1.5. Question Time and Response Time
The study was designed for patients and attendants to use the bot seven days pre-surgery and seven days post-surgery. However, due to a bug in the code, the bot was never disabled. Moreover, 3 participants had their surgeries rescheduled, leading them to use the bot for a longer period. Overall, participants sent messages to the bot from 19 days before their surgeries to 24 days post their surgeries (Figure 4B). A majority of messages (71.4%) were sent in the 8-day period from 5 days prior to surgery to 2 days post-surgery with the highest number of messages sent on the day of the surgery (76). 156 messages were sent before the surgery day and 111 post the surgery day. We conducted a one-way ANOVA on the number of messages asked by each patient-attendant group pre-surgery, on the day of surgery, and post-surgery, and found statistical significance with F(2, 60)=3.69, p¡.05. Post-hoc pairwise comparisons with Bonferroni correction indicated a statistically significant difference between ‘pre’ and ‘on’ surgery (t(30)=2.7, p¡.05) and between ‘pre’ and ‘post’ surgery (t(30)=3.2, p¡.05). This suggests that patients and attendants are more inquisitive before surgery, potentially due to heightened anxiety, aligning with prior research (Beisecker and Beisecker, 1990). The higher pre-surgery usage may also be due to the novelty effect. With respect to the types of queries, logistical questions were mainly asked before the surgery, reaching their peak one day before the surgery day (Figure 4B), as anticipated. In contrast, experts expressed that questions on the day of surgery might be unnecessary:
“Multiple questions were asked on the day of surgery, “when he is going to be discharged?” … You are in the hospital… Why to use the bot!… You can walk 10 steps and go to the staff and ask.” – D1.
Regarding the usage patterns of experts, responses were verified throughout the day. The only exception was during days reserved for surgeries. All doctors emphasized a no-phone-use policy during surgeries, stressing the need for minimal messages on those days. A coordinator suggested adding a working-time parameter for CataractBot to restrict message delivery to specified periods. Experts refrained from using their phones in front of patients due to the negative perception associated with using phones for personal reasons. During working hours, the long queue of waiting patients made it challenging for experts to use their phones. These contributed to several queries being escalated. Opinions on using the chatbot after work hours varied: two experts were reluctant, while four were in favor of using it “anytime”.
@@ 6.2. Features
6.2.1. Experts-in-the-loop
@@
The novel feature of having human experts verify and update the auto-generated responses was highly appreciated by participants, serving as the key reason for their engagement with the bot. For instance, “That’s why I gave my WhatsApp (number to onboard the bot), because I can talk directly to the doctor.” – A2. Contrary to initial expectations that understanding the icons might be complex, participants found it self-explanatory:
“That’s simple… Even my dad understood. I didn’t tell him the tick mark, question mark… He told me, “okay, doctor verified… Now the tick mark came”… If MY dad was able to use, I think anyone can.” – A13.
@@
However, since participants knew that these messages were being reviewed by doctors and coordinators, four used the bot to send direct messages to them instead of asking questions. For instance, P19 messaged “Dr. Give me a call?”
The length of expert corrections in characters varied significantly (t(91)=5.2, p¡0.05) with 152.4104.9 for medical questions and 50.942.6 for logistical questions. Further analysis revealed that their suggested edits fell into five categories: adding missing information (76.8%), asking clarification question (14.6%), factual corrections (11.0%), asking the patient to visit the hospital (8.8%), and removing extra unnecessary information (3.7%).
@@
Finally, experts did not have access to the final updated answer sent to information seekers after their suggested edits. This design decision limited verification to a single round, minimizing the workload of experts. However, without this transparency, experts were sometimes unsure whether or how CataractBot had understood their corrections, especially when the edits aimed at improving readability and structure rather than solely focusing on accuracy and completeness: “I’d asked to flip [the answer] around… to give information for local anesthesia first then for general anesthesia… Did the bot actually understood and give that as an answer? I donno.” (D3). In the CataractBot system, only the KB update expert had access to the updated answer sent to the information seeker.